OpenStudy (anonymous):

Y-3=-3x+2 in standard form

OpenStudy (texaschic101):

y - 3 = -3x + 2 y = -3x + 2 +3 y = -3x + 5 -3x + y = 5 (multiply by -1 to make x positive) 3x - y = -5 <== standard form

OpenStudy (texaschic101):

any questions ?